OpenStudy (anonymous):

Jake's owners are going on vacation and they are going to board him at a dog kennel. The Happy Puppies Kennel charges $15.00 per night. The Mutt Hutt Kennel charges a flat fee of $75.00 plus $7.00 per night. If Jake will be staying at the kennel for 12 nights, which kennel will be the least expensive? A. There is not enough information to decide. B. They are both the same price. C. The Happy Puppies Kennel would be the least expensive. D. The Mutt Hutt Kennel would be the least expensive.

OpenStudy (texaschic101):

Happy Puppies kennel -- 15(12) = 180 Mutt Hutt -- 75 + 7(12) = 75 + 84 = 159 I agree...